B. If you brake hard and fast, you can bring a typical car to rest from 100 kph (about 60 mph) in 5 seconds.

B.1 Calculate the magnitude of your acceleration, −a0, assuming that it is constant. Answer is in m/s2
Answer #1

Initially velocity u=100 kph = 100*5/18 = 27.77 m/s

Final velocity v= 0 m/s

Acceleration a =\frac{v-u}{t}=\frac{0-27.77}{5} =-5.56 m/s^2